# Shuofei Qiao♠♡, Pengjun Xie, Fei Huang, Huajun Chen, Ningyu Zhang

### v1

@@ Introduction
As large language models (LLMs) grow ever more powerful, LLM-based agents augmented by their own reasoning and external tools are taking on increasingly sophisticated works (Zhao et al. 2023; Wang et al. 2024a; Xi et al. 2025; Qiao et al. 2023). No longer mere assistants, these agents now trawl the web for elusive insights and weave them into comprehensive, publication ready reports, like Deep Research (OpenAI 2025; x.ai 2025) and WebDancer (Wu et al. 2025a). Moreover, they can handle complex data analyses (Lan et al. 2025; Ifargan et al. 2025; Ou et al. 2025), navigate multi-step GUI workflows (Luo et al. 2025; Qin et al. 2025), and sustain long-horizon, tool-rich interactions (Yao et al. 2025; Barres et al. 2025; Chen et al. 2025; Fang et al. 2025) with precision. Yet executing such intricate, long-horizon tasks demands dozens of steps and protracted runtimes. Along the way, unpredictable external events—network glitches, UI changes, shifting data schemas—can derail the entire process. Restarting from scratch every time is a punishing ordeal for present-day agents. Beneath their surface diversity, many complex tasks share deep structural commonalities and a similar environment. Instead of starting fresh each time, an agent should extract its experience from past successes. By turning earlier trajectories into reusable templates like patterns of reasoning, tool sequences, and recovery tactics, it can progress step by step, learning from every failure and success, until even the most convoluted missions become routine.
@@
The capacity to distill, chronicle, and re-apply lessons from one’s own experiential trajectory is the bedrock of human learning and the pivotal gateway through which an agent ascends toward self-directed refinement (Liu et al. 2025a; Sumers et al. 2023; Li et al. 2023). Procedural memory (Gupta and Cohen 2002; Cohen and Squire 1980) silently compiles habitual skills into executable subroutines, enabling unconscious, fluent action. While contemporary agents built on LLMs can compose short action plans or call external tools, their procedural knowledge is either hand-crafted, stored as brittle prompt templates, or implicitly entangled in model parameters that are expensive to update. Existing memory-augmented frameworks such as LangGraph (Mavroudis 2024), AutoGPT (Yang, Yue, and He 2023), or agent cognitive architectures like Memory Bank (Zhong et al. 2024a) and Soar (Laird 2022) provide coarse abstractions (buffers, rule chunks, production systems) but leave the optimization of procedural memory life-cycle operations about how skills are built, indexed, patched, and eventually pruned, largely unexamined. Consequently, there is no principled way to quantify how efficiently an agent evolves its procedural repertoire or to guarantee that new experiences improve rather than erode performance.
@@
Memory is a foundational component in language agents, enabling them to retain and utilize past information across multiple timescales, including short-term, episodic, and long-term memory, to enhance their performance and adaptability (Zhou et al. 2023, 2024; Zhang et al. 2024; Liu et al. 2025a). These systems aim to mimic aspects of human memory to improve coherence, personalization, and learning capabilities  (Chhikara et al. 2025; Wu et al. 2025b). Current approaches include end-to-end memory systems (Yu et al. 2025; Zhou et al. 2025), external memory systems (Chhikara et al. 2025; Zhong et al. 2024b), and hierarchical memory structures (Hu et al. 2024a; Xu et al. 2025). These methods involve encoding and storing information in various formats, using retrieval mechanisms like vector embeddings and semantic search, and implementing memory updating and forgetting strategies to maintain relevance and efficiency. Despite its importance, memory in multi-turn agent interactions remains underexplored, and enabling agents to effectively learn and utilize memory across trajectories poses a significant challenge. Procedural memory is a type of long-term memory that involves the retention of procedures and skills, such as typing or riding a bike, which are performed automatically without conscious thought. The agent utilizes procedural memory to internalize and automate repetitive tasks, decision-making processes, and interaction patterns, leading to more efficient and context-aware responses over time. Although there have been several works, such as Voyager (Wang et al. 2023), AWM (Wang et al. 2024b), and AutoManual (Chen et al. 2024), that utilize procedural memory to enhance agents’ capabilities on similar tasks, there still lacks a systematic analysis on how to construct, retrieve, and update such procedural memory. Therefore, our work mainly focuses on exploring how to build an effective procedural memory system for agents performing cross-trajectory tasks.
@@
LLM-based Agent learning from experience involves intelligence continuously improving their decision-making capabilities through interaction with environments and utilization of past experiences (Tan et al. 2025; Tang et al. 2025; Zhou et al. 2025; Qiao et al. 2025). This approach is crucial for developing adaptive and intelligent agents capable of handling dynamic real-world scenarios, as it allows them to optimize behaviors, reduce manual programming needs, and enhance performance across various tasks (Zheng et al. ; Liu et al. 2025c; Wang et al. 2025). Agents typically employ mechanisms such as reinforcement learning (Lu et al. 2025; Dong et al. 2025), experience replay (Feng et al. 2025; Liu et al. 2025b), imitation learning (Sun et al. 2024; Yang et al. 2024b), memory management (Hou, Tamoto, and Miyashita 2024; Hu et al. 2024b), and multi-agent learning to achieve this. However, current methods face limitations including low sample efficiency, poor generalization across tasks, catastrophic forgetting when learning new information, and there are very few features for memory update. Additionally, collecting high-quality training data can be challenging and may introduce biases. Addressing these limitations is essential for advancing the capabilities of LLM-based agents and ensuring their effective application in real-world contexts.

### v2

@@ Introduction
As large language models (LLMs) grow ever more powerful, LLM-based agents augmented by their own reasoning and external tools are taking on increasingly sophisticated works (Zhao et al. 2023; Wang et al. 2024a; Xi et al. 2025; Qiao et al. 2023). No longer mere assistants, these agents now trawl the web for elusive insights and weave them into comprehensive, publication ready reports, like Deep Research (OpenAI 2025; x.ai 2025) and WebDancer (Wu et al. 2025a). Moreover, they can handle complex data analyses (Lan et al. 2025; Ifargan et al. 2025; Ou et al. 2025), navigate multi-step GUI workflows (Luo et al. 2025; Qin et al. 2025), and sustain long-horizon, tool-rich interactions (Yao et al. 2025; Barres et al. 2025; Chen et al. 2025; Fang et al. 2025; Gur et al. 2023) with precision. Yet executing such intricate, long-horizon tasks demands dozens of steps and protracted runtimes. Along the way, unpredictable external events—network glitches, UI changes, shifting data schemas—can derail the entire process. Restarting from scratch every time is a punishing ordeal for present-day agents. Beneath their surface diversity, many complex tasks share deep structural commonalities and a similar environment. Instead of starting fresh each time, an agent should extract its experience from past successes. By turning earlier trajectories into reusable templates like patterns of reasoning, tool sequences, and recovery tactics, it can progress step by step, learning from every failure and success, until even the most convoluted missions become routine.
@@
The capacity to distill, chronicle, and re-apply lessons from one’s own experiential trajectory is the bedrock of human learning and the pivotal gateway through which an agent ascends toward self-directed refinement (Liu et al. 2025a; Sumers et al. 2023a; Li et al. 2023). Procedural memory (Gupta and Cohen 2002; Cohen and Squire 1980) silently compiles habitual skills into executable subroutines, enabling unconscious, fluent action. While contemporary agents built on LLMs can compose short action plans or call external tools, their procedural knowledge is either hand-crafted, stored as brittle prompt templates, or implicitly entangled in model parameters that are expensive to update. Existing memory-augmented frameworks such as LangGraph (Mavroudis 2024), AutoGPT (Yang, Yue, and He 2023), or agent cognitive architectures like Memory Bank (Zhong et al. 2024a; Sumers et al. 2023b) and Soar (Laird 2022) provide coarse abstractions (buffers, rule chunks, production systems) but leave the optimization of procedural memory life-cycle operations about how skills are built, indexed, patched, and eventually pruned, largely unexamined. Consequently, there is no principled way to quantify how efficiently an agent evolves its procedural repertoire or to guarantee that new experiences improve rather than erode performance.
@@
Memory is a foundational component in language agents, enabling them to retain and utilize past information across multiple timescales, including short-term, episodic, and long-term memory, to enhance their performance and adaptability (Zhou et al. 2023, 2024; Zhang et al. 2024; Liu et al. 2025a; Li et al. 2025). These systems aim to mimic aspects of human memory to improve coherence, personalization, and learning capabilities  (Chhikara et al. 2025; Wu et al. 2025b; Xia et al. 2025). Current approaches include end-to-end memory systems (Yu et al. 2025; Zhou et al. 2025), external memory systems (Chhikara et al. 2025; Zhong et al. 2024b), and hierarchical memory structures (Hu et al. 2024a; Xu et al. 2025). These methods involve encoding and storing information in various formats, using retrieval mechanisms like vector embeddings and semantic search, and implementing memory updating and forgetting strategies to maintain relevance and efficiency. Despite its importance, memory in multi-turn agent interactions remains underexplored, and enabling agents to effectively learn and utilize memory across trajectories poses a significant challenge. Procedural memory is a type of long-term memory that involves the retention of procedures and skills, such as typing or riding a bike, which are performed automatically without conscious thought. The agent utilizes procedural memory to internalize and automate repetitive tasks, decision-making processes, and interaction patterns, leading to more efficient and context-aware responses over time. Although there have been several works, such as Voyager (Wang et al. 2023), AWM (Wang et al. 2024b), and AutoManual (Chen et al. 2024), that utilize procedural memory to enhance agents’ capabilities on similar tasks, there still lacks a systematic analysis on how to construct, retrieve, and update such procedural memory like  (Wu et al. 2024). Therefore, our work mainly focuses on exploring how to build an effective procedural memory system for agents performing cross-trajectory tasks.
@@
LLM-based Agent learning from experience involves intelligence continuously improving their decision-making capabilities through interaction with environments and utilization of past experiences (Tan et al. 2025; Tang et al. 2025; Zhou et al. 2025; Qiao et al. 2025; Su et al. 2025; Wang et al. 2024b). This approach is crucial for developing adaptive and intelligent agents capable of handling dynamic real-world scenarios, as it allows them to optimize behaviors, reduce manual programming needs, and enhance performance across various tasks (Zheng et al. ; Liu et al. 2025c; Wang et al. 2025). Agents typically employ mechanisms such as reinforcement learning (Lu et al. 2025; Dong et al. 2025), experience replay (Feng et al. 2025; Liu et al. 2025b), imitation learning (Sun et al. 2024; Yang et al. 2024b), memory management (Hou, Tamoto, and Miyashita 2024; Hu et al. 2024b), and multi-agent learning to achieve this. However, current methods face limitations including low sample efficiency, poor generalization across tasks, catastrophic forgetting when learning new information, and there are very few features for memory update. The key distinction of our work lies in systematically investigating optimal strategies for construction, retrieval, and update modules of an agent’s procedural knowledge. During the update phase, we enhance the agent’s capabilities by maintaining an editable repository of procedural knowledge. Additionally, collecting high-quality training data can be challenging and may introduce biases. Addressing these limitations is essential for advancing the capabilities of LLM-based agents and ensuring their effective application in real-world contexts.